OpenStudy (anonymous):

6= sin( (2pi x/13) + pi/2)) +1.7 Solve for x?

OpenStudy (anonymous):

impossible

OpenStudy (anonymous):

ok, the question is: The depth of water, d metres at a port entrance is modelled by the function: d(t)=sin( (2pi x/13) + pi/2)) +1.7 (t= time in hrs since 12 noon) The ship needs 1.2 m of water to sail into or out of port. She needs two hrs to get from the entrance of the port to the jetty; 1 1/2 hrs to unload and 2 hours to depart the jetty and sail back to the port entrance. Use exact values to calculate the earliest time HMAS King is able to complete its activities and be back at the port entrance?

OpenStudy (anonymous):

1)2piX/13+pi/2=arc sin (4.3)+2n(pi)------solve X,its easy 2)2piX/13+pi/2=-arc sin(4.3)+(2n+1)pi---------solve X..there r two solution...check it yourself to get the range

OpenStudy (anonymous):

but its not using exact values?

OpenStudy (anonymous):

oh i dont have calculator...sry for that...

OpenStudy (anonymous):

umm i meant, like using pi/6 pi/3 etc

OpenStudy (anonymous):

put pi=3.14,since its in radian..

OpenStudy (anonymous):

i meant like use the unit circle to solve for x

OpenStudy (anonymous):

oh buddy u just got me confused here by putting the equations value=6 earlier which is wrong coz arc sin wont give any value since its range is -1 to+1....ok to solve this problem u need to differentiate the d(x) and put d'(x)=0 and find x(or 't' in this case),see how many points u get,those r critical points,put these values in the equation and check which gives the minimum time and that will be ur solution